OpenStudy (anikate):

Commutative/ Associative Property of Addition, Commutative/ Associative property of Multiplication, Identity property of Addition, Identity property of multiplication, zero product property ? please explain all these to me?

OpenStudy (agent47):

That is a lot.... Addition: commutative: a+b=b+a associative: (a+b)+c = a+(b+c)

OpenStudy (anikate):

ya i know its a lot

OpenStudy (agent47):

Multiplication: commutative: a*b=b*a associative: a(bc)=(ab)c

OpenStudy (agent47):

Identity property of addition: a+0=a (not sure abt the identity property of addition) identity property of multiplication: a*1=a (sure abt this one)

OpenStudy (agent47):

zero product property - idk, but i think it means a*0=0

OpenStudy (agent47):

Google does a better job than me

OpenStudy (anikate):

no you do better

OpenStudy (anikate):

-2+(1+x)=(-2+1)+x what property? @Agent47

OpenStudy (anikate):

@Agent47

OpenStudy (anikate):

@Agent47

OpenStudy (anikate):

@Agent47

OpenStudy (anikate):

@Agent47 help me plz

OpenStudy (anikate):

-2+(1+x)=(-2+1)+x what property? @Agent47

OpenStudy (anikate):

i need help on 2 others after this

OpenStudy (anonymous):

associative property of addition, as the parentheses have moved but the symbols stayed in the same order